the vapor pressure of H2O at 28 degrees Celsius is 28.3mmhg find the vapor pressure of a solution containing 100g of sugar c12H22011 in 500g of H20 at 28 degrees celcius.

1
Expert's answer
2022-03-22T09:44:03-0400

M(sugar, C12H22O11) = 342 g/mole;

m(sugar, C12H22O11) = 100 g;

M(H2O) = 18 g/mole;

m(H2O) = 500 g;

P0 = 28.3mmHg;

n(sugar, C12H22O11) = m(sugar, C12H22O11)/M(sugar, C12H22O11) = 100/342 = 0.29 moles;

n(H2O) = m(H2O)/M(H2O) = 500/18 = 27.78 moles;

xi = n(H2O)/(n(H2O) + n(sugar, C12H22O11)) = 27.78/(27.78 + 0.29) = 27.78/28.07 = 0.9897

Psolution = xi * P0 = 0.9897 * 28.3 = 28.0 mmHg.

Answer: 28.0 mmHg.